# Do we need domain-specific embedding models? An empirical investigation

### v3

@@ Abstract
Embedding models play a crucial role in representing and retrieving information across various NLP applications. Recent advancements in Large Language Models (LLMs) have further enhanced the performance of embedding models, which are trained on massive amounts of text covering almost every domain. These models are often benchmarked on general-purpose datasets like Massive Text Embedding Benchmark (MTEB), where they demonstrate superior performance. However, a critical question arises: Is the development of domain-specific embedding models necessary when general-purpose models are trained on vast corpora that already include specialized domain texts? In this paper, we empirically investigate this question, choosing the finance domain as an example. We introduce the Finance Massive Text Embedding Benchmark (FinMTEB), a counterpart to MTEB that consists of financial domain-specific text datasets. We evaluate the performance of seven state-of-the-art embedding models on FinMTEB and observe a significant performance drop compared to their performance on MTEB. To account for the possibility that this drop is driven by FinMTEB’s higher complexity, we propose four measures to quantify dataset complexity and control for this factor in our analysis. Our analysis provides compelling evidence that state-of-the-art embedding models struggle to capture domain-specific linguistic and semantic patterns. Moreover, we find that the performance of general-purpose embedding models on MTEB is not correlated with their performance on FinMTEB, indicating the need for domain-specific embedding benchmarks for domain-specific embedding models. This study sheds light on developing domain-specific embedding models in the LLM era. FinMTEB comes with open-source code at https://github.com/yixuantt/FinMTEB
@@ 3.1 FinMTEB Tasks
Figure 1: An overview of tasks and datasets used in FinMTEB. All the dataset descriptions and examples are provided in the Appendix C.
@@
In summary, FinMTEB consists of a total of 64 datasets, spanning seven different tasks. The key difference between MTEB and FinMTEB is that all datasets in FinMTEB are finance-domain specific, either previously used in financial NLP research or newly developed by the authors. Semantic similarity between datasets in FinMTEB are shown in Appendix A. The detailed dataset information and descriptions are presented in the Appendix C. The main scoring metric for each task is the same as used with that of the MTEB benchmark, and the details are presented in the Appendix D.
@@ 3.2 Characteristics of FinMTEB
High Semantic Diversity. We first examine the inter-dataset semantic similarity. Using the all-MiniLM-L6-v2 model [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, we embed 100 samples from each dataset, compute their averages to represent the dataset embedding, and measure inter-dataset similarity using cosine similarity. As shown in Appendix A, most datasets in FinMTEB have an inter-dataset similarity score below 0.6, indicating high semantic diversity. Despite being finance-domain specific, this highlights the diverse narratives and contexts present in the financial datasets.
@@ 3.3 The performance of state-of-the-art embedding models on FinMTEB
Table 2: State-of-the-art Embedding Model Performance on MTEB and FinMTEB. The “MTEB Score” column represents performance on the MTEB benchmark (Muennighoff et al., 2022) as reported on the Hugging Face MTEB Leaderboard  [^5] https://huggingface.co/spaces/mteb/leaderboard. The “FinMTEB Score” column shows the average performance score evaluated on the proposed FinMTEB benchmark. To ensure a fair comparison, FinMTEB uses the same evaluation metrics as MTEB.
@@ 3.3.1 What Drives the Performance Gap?
We present the ANOVA analysis results in Appendix B. First, the results indicate that the choice of embedding model (Model factor) significantly impacts performance in most tasks, such as STS (0.79), Retrieval (0.89), and Reranking (0.30), with the exception of Pair-classification (0.00), where model choice has no significant impact. Second, the Domain factor also shows significant effects across all embedding tasks. Interestingly, the average scores reveal that, from an overall perspective, the Model factor has little impact on performance, with an effect size of 0.00 and an insignificant p-value. This suggests that while individual models may excel at specific tasks, their performance discrepancies balance out when averaged. However, the Domain factor (1.00) demonstrates a much more prominent influence, underscoring the necessity for domain-specific models or fine-tuning when addressing specialized tasks like those in finance.
@@ 4.1 Quantifying Dataset Complexity
ChatGPT Error Rate. The first measure quantifies how challenging it is for ChatGPT to answer a dataset’s questions. Specifically, for each example in the dataset across different tasks, we reformat the example into a question-and-answer pair, as shown in Appendix F, and prompt GPT-4o mini (OpenAI, 2024a). The rationale is that if ChatGPT fails to answer a question correctly, it indicates the difficulty level of the question. Additionally, since state-of-the-art LLM-based embedding models present each query with an instruction in a question-answer format, we use the ChatGPT error rate as an indicator of dataset complexity.
@@
For all of these four complexity measures, a higher score indicates higher dataset complexity. Details on the measures and their calculations are provided in Appendix E.
@@ 5 Domain-specific Embedding Benchmark is needed
Table 5: Spearman’s correlation of embedding models’ performance on MTEB and FinMTEB across different tasks. The p-value indicates that all correlations are statistically insignificant, suggesting no relationship between embedding model performance on the two benchmarks.
@@ Appendix A Dataset Semantic Similarity
Appendix B ANOVA DATA
@@
Appendix C Datasets
The detailed description of each dataset used in this work is listed in the Table tables 9, 12, 10, 11, 7, 13 and 8.
@@ Appendix C Datasets
Table 7: Summary of STS Datasets
@@
Table 8: Summary of Retrieval Datasets
@@
Table 9: Summary of Classification Datasets
@@
Table 10: Summary of Clustering Datasets
@@
Table 11: Summary of Summarization Datasets
@@
Table 12: Summary of Reranking Datasets
@@
Table 13: Summary of PairClassification Datasets
Appendix D Main Metric
@@
Appendix E Complexity Score Calculation
E.1 Shannon Entropy in Information Theory
@@
E.2 Readability: Gunning Fog Index
@@
E.3 Mean Dependency Distance
@@
Appendix F Prompt For ChatGPT Error Rate
The detailed example prompt of each task for the ChatGPT Error Rate is listed in Table tables 14, 17, 16, 20, 18, 19 and 15.
@@ Appendix F Prompt For ChatGPT Error Rate
Table 15: Prompt for the ChatGPT Error Rate on Classification.
@@
Table 16: Prompt for the ChatGPT Error Rate on Retrieval.
@@
Table 17: Prompt for the ChatGPT Error Rate on Clustering.
@@
Table 18: Prompt for the ChatGPT Error Rate on Reranking.
@@
Table 19: Prompt for the ChatGPT Error Rate on Pair-Classification.
@@
Table 20: Prompt for the ChatGPT Error Rate on Summarization.

### v4

@@ Abstract
Embedding models play a crucial role in representing and retrieving information across various NLP applications. Recent advancements in Large Language Models (LLMs) have further enhanced the performance of embedding models, which are trained on massive amounts of text covering almost every domain. These models are often benchmarked on general-purpose datasets like Massive Text Embedding Benchmark (MTEB), where they demonstrate superior performance. However, a critical question arises: Is the development of domain-specific embedding models necessary when general-purpose models are trained on vast corpora that already include specialized domain texts? In this paper, we empirically investigate this question, choosing the finance domain as an example. We introduce the Finance Massive Text Embedding Benchmark (FinMTEB), a counterpart to MTEB that consists of financial domain-specific text datasets. We evaluate the performance of seven state-of-the-art embedding models on FinMTEB and observe a significant performance drop compared to their performance on MTEB. To account for the possibility that this drop is driven by FinMTEB’s higher complexity, we propose four measures to quantify dataset complexity and control for this factor in our analysis. Our analysis provides compelling evidence that state-of-the-art embedding models struggle to capture domain-specific linguistic and semantic patterns. Moreover, we find that the performance of general-purpose embedding models on MTEB is not correlated with their performance on FinMTEB, indicating the need for domain-specific embedding benchmarks for domain-specific embedding models. This study sheds light on developing domain-specific embedding models in the LLM era.
@@ 3.1 FinMTEB Tasks
Figure 1: An overview of tasks and datasets used in FinMTEB. All the dataset descriptions and examples are provided in the Appendix D.
@@
In summary, FinMTEB consists of a total of 64 datasets, spanning seven different tasks. The key difference between MTEB and FinMTEB is that all datasets in FinMTEB are finance-domain specific, either previously used in financial NLP research or newly developed by the authors. Semantic similarity between datasets in FinMTEB are shown in Appendix A. The detailed dataset information and descriptions are presented in the Appendix D. The main scoring metric for each task is the same as used with that of the MTEB benchmark, and the details are presented in the Appendix E.
@@ 3.2 Characteristics of FinMTEB
Semantic Diversity. We examine the inter-dataset semantic similarity. Using the all-MiniLM-L6-v2 model [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, we embed 1000 samples from each dataset, compute their averages to represent the dataset embedding, and measure inter-dataset similarity using cosine similarity. As shown in Appendix A, most datasets in FinMTEB have an inter-dataset similarity score below 0.6, with a mean cosine similarity of 0.4. Despite being finance-domain specific, this highlights the diverse narratives and contexts present in the financial datasets.
@@ 3.3 The performance of state-of-the-art embedding models on FinMTEB
Table 2: State-of-the-art Embedding Model Performance on MTEB and FinMTEB. The “MTEB Score” column represents performance on the MTEB benchmark (Muennighoff et al., 2022) as reported on the Hugging Face MTEB Leaderboard  [^5] https://huggingface.co/spaces/mteb/leaderboard. The “FinMTEB Score” column shows the average performance score evaluated on the proposed FinMTEB benchmark. To ensure a fair comparison, FinMTEB uses the same evaluation metrics as MTEB. More evaluation results on other SoTA models are presented in Appendix B.
@@ 3.3.1 What Drives the Performance Gap?
We present the ANOVA analysis results in Appendix C. First, the results indicate that the choice of embedding model (Model factor) significantly impacts performance in most tasks, such as STS (0.79), Retrieval (0.89), and Reranking (0.30), with the exception of Pair-classification (0.00), where model choice has no significant impact. Second, the Domain factor also shows significant effects across all embedding tasks. Interestingly, the average scores reveal that, from an overall perspective, the Model factor has little impact on performance, with an effect size of 0.00 and an insignificant p-value. This suggests that while individual models may excel at specific tasks, their performance discrepancies balance out when averaged. However, the Domain factor (1.00) demonstrates a much more prominent influence, underscoring the necessity for domain-specific models or fine-tuning when addressing specialized tasks like those in finance.
@@ 4.1 Quantifying Dataset Complexity
ChatGPT Error Rate. The first measure quantifies how challenging it is for ChatGPT to answer a dataset’s questions. Specifically, for each example in the dataset across different tasks, we reformat the example into a question-and-answer pair, as shown in Appendix G, and prompt GPT-4o mini (OpenAI, 2024a). The rationale is that if ChatGPT fails to answer a question correctly, it indicates the difficulty level of the question. Additionally, since state-of-the-art LLM-based embedding models present each query with an instruction in a question-answer format, we use the ChatGPT error rate as an indicator of dataset complexity.
@@
For all of these four complexity measures, a higher score indicates higher dataset complexity. Details on the measures and their calculations are provided in Appendix F.
@@ 5 Domain-specific Embedding Benchmark is needed
Table 5: Spearman’s correlation of embedding models’ performance on MTEB and FinMTEB across different tasks. The p-value indicates that all correlations are statistically insignificant, suggesting a lack of evidence for a relationship between embedding model performance on the two benchmarks.
@@ Appendix A Dataset Semantic Similarity
Appendix B Additional State-of-the-art Embedding Model Performance
Table 6: Performance metrics of more state-of-the-art (SoTA) models on FinMTEB and their testing times (batch size = 512)
Model STS Retr. Rerank. Class. Summ. PairClass. Clus. Avg. Duration (/hr)
Echo Embedding (Springer et al., 2024) 0.4380 0.6443 0.9765 0.6525 0.4722 0.6261 0.5776 0.6267 12.00
AnglE-BERT (Li & Li, 2023) 0.3080 0.5730 0.9650 0.6439 0.5049 0.6891 0.5774 0.6088 8.00
NV-Embed v2 (Lee et al., 2024) 0.3739 0.7061 0.9822 0.6393 0.5103 0.6043 0.6096 0.5739 5.60
BeLLM (Li & Li, 2024) 0.3919 0.0169 0.5661 0.1168 0.3906 0.5682 0.0685 0.3027 11.98
Appendix C ANOVA DATA
@@
Appendix D Datasets
The detailed description of each dataset used in this work is listed in the Table tables 10, 13, 11, 12, 8, 14 and 9.
Table 8: Summary of STS Datasets
@@ Appendix D Datasets
Table 9: Summary of Retrieval Datasets
@@
Table 10: Summary of Classification Datasets
@@
Table 11: Summary of Clustering Datasets
@@
Table 12: Summary of Summarization Datasets
@@
Table 13: Summary of Reranking Datasets
@@
Table 14: Summary of PairClassification Datasets
@@
Appendix E Main Metric
@@
Appendix F Complexity Score Calculation
F.1 Shannon Entropy in Information Theory
@@
F.2 Readability: Gunning Fog Index
@@
F.3 Mean Dependency Distance
@@
Appendix G Prompt For ChatGPT Error Rate
The detailed example prompt of each task for the ChatGPT Error Rate is listed in Table tables 15, 18, 17, 21, 19, 20 and 16.
@@ Appendix G Prompt For ChatGPT Error Rate
Table 16: Prompt for the ChatGPT Error Rate on Classification.
@@
Table 17: Prompt for the ChatGPT Error Rate on Retrieval.
@@
Table 18: Prompt for the ChatGPT Error Rate on Clustering.
@@
Table 19: Prompt for the ChatGPT Error Rate on Reranking.
@@
Table 20: Prompt for the ChatGPT Error Rate on Pair-Classification.
@@
Table 21: Prompt for the ChatGPT Error Rate on Summarization.
@@
Appendix H Supplementary Experiment: Finetune SoTA Embedding Using Domain Corpus
We fine-tuned e5-mistral-7b-instruct (Wang et al., 2023) using a syntenic finance QA dataset generated through Self-instruct(Wang et al., 2022) with GPT-4o mini (OpenAI, 2024a) and a manually labeled seed finance dataset. The results demonstrate clear domain adaptation effects:
On FinMTEB (Table 22), performance improved from 0.6475 to 0.6735, showing the benefits of finance-specific training. While general MTEB scores (Table 23) slightly decreased from 0.6463 to 0.6320, the model maintained competitive performance on broader tasks.
These results highlight how domain adaptation can enhance specialized task performance while preserving general capabilities.
Table 22: Overall Performance on FinMTEB
Model STS Ret. Rerank. Class. Summ. PairClass. Clust. Avg.
e5-mistral-7b-instruct 0.380 0.675 0.988 0.645 0.528 0.739 0.578 0.648
   + domain adaption 0.428 0.699 0.990 0.757 0.480 0.801 0.560 0.674
Table 23: Overall Performance on MTEB
Model STS Ret. Rerank. Class. Summ. PairClass. Clust. Avg.
e5-mistral-7b-instruct 0.859 0.588 0.602 0.770 0.314 0.883 0.508 0.646
   + domain adaption 0.858 0.491 0.603 0.776 0.306 0.875 0.517 0.632